OpenStudy (anonymous):

Find the area of the region Y=x*sqrt((4-x)/(4+x)), y=0, x=4

OpenStudy (jdoe0001):

just plug in the values :) given for "x" and "y" in the equation

OpenStudy (anonymous):

It's integration

OpenStudy (anonymous):

\[y=x\sqrt{\frac{4-x}{4+x}}\] Here's a link with a pic of the region: http://www.wolframalpha.com/input/?i=Plot%5Bx*Sqrt%5B%284-x%29%2F%284%2Bx%29%5D%2C%7Bx%2C-1%2C4%7D%5D \[A=\int_0^4 x\sqrt{\frac{4-x}{4+x}}~dx\] I'd start off with a substitution: \(u=4+x~\iff~x=u-4\), so \(du=dx\): \[A=\int_4^8 (u-4)\sqrt{\frac{4-(u-4)}{u}}~du\\ A=\int_4^8 (u-4)\sqrt{\frac{8-u}{u}}~du \] Then another substitution: \(t=\dfrac{8-u}{u}~\iff~u=\dfrac{8}{t+1}\), so \(dt=-\dfrac{8}{u^2}~du~\iff~du=-\dfrac{\left(\frac{8}{t+1}\right)^2}{8}~dt=-\dfrac{8}{(t+1)^2}~dt\). So you have \(u-4=\dfrac{8}{t+1}-4=\dfrac{4-4t}{t+1}\), and the integral becomes \[A=\int_1^0 \dfrac{4-4t}{t+1}\sqrt{t}~\left(-\dfrac{8}{(t+1)^2}~dt\right)\\ A=32\int_0^1 \dfrac{1-t}{(t+1)^3}\sqrt{t}~dt\] Yet another substitution: \(s=\sqrt{t}~\iff~s^2=t\), so \(ds=\dfrac{1}{2\sqrt{t}}~dt~\iff~dt=2s~ds\). \[A=32\int_0^1 \dfrac{1-s^2}{(s^2+1)^3}s~(2s~ds)\\ A=64\int_0^1 \dfrac{(1-s^2)s^2}{(s^2+1)^3}~ds\] Break up into partial fractions. The work from there gets easier.